OpenStudy (anonymous):

Solve the following system of equations. 4x + 7y – 2z = 0 3x – 5y + 3z = 9 3x + 6y – z = 1 (–1, 0, –2) (1, 0, –2) (–1, 0, 2) (1, 0, 2)

OpenStudy (anonymous):

(1, 0, 2)

OpenStudy (anonymous):

when they give you multiple solutions just pick one and susbtitute the values on the equation. If it fits, then you have your answer for example: 4x + 7y – 2z = 0 , (1, 0, 2) 4(1) +7(0) -2(2) = 0 4-4 = 0 check

jhonyy9 (jhonyy9):

the 4th